Which of the following scenario is correct?

1. Sales price with credit term = a (for example) sales tax = 5% sales return = b Credit term discount = c (I am not sure here if it is applied after tax or before tax) payment made by customer? I suppose that Initial Sales = a + 0.05a = 1.05a Sales return = b + 0.05b = 1.05b So the actual sales = 1.05a - 1.05b After discount/ Final sales= (1.05a - 1.05b) - c

2. All detail is same as above except calculation Initial Sale = a Sale return = b Actual sale = a - b After disount = (a-b)-c After tax/ Actual sales = 1.05 ((a-b)-c)

3. All detail as above Initial Sale = 1.05a Sale return = 1.05b discount applied = (a-b)-c After tax / Actual tax = 1.05 ((a-b)-c)

And my final concern is if scenario 1 is correct then my question is how can a discount be applied on tax? if scenario 2 is correct then how can a company sell their product without sale tax? and if scenario 3 is correct then what about tax liability (as a company view point), by the time sales return is made company may already paid the tax to state or journalize/posted on ledger; how to accout it later?
Answer #1

Scenario 2 is correct

2. All detail is same as above except calculation Initial Sale = a Sale return = b Actual sale = a - b After discount = (a-b)-c After tax/ Actual sales = 1.05 ((a-b)-c)

because the tax or sales tax applies to the final value derived after sales return and all discounts. The value derived after will be returns and discount will be (a-b-c) and then 5% sales tax will be applied to it.